How do you work this equation out? You have to find x and y
(2x-3i)+(-3+2y)i=y-xi

Answer by stanbon(75887)

find x and y
(2x-3i)+(-3+2y)i=y-xi
-----
Comment: Complex numbers have a Real part and
an Imaginary part.
Real part Equation::
2x - 3 = y
Imaginary part Equation::
-3 + 2y = -x
-----
Rearrange:
2x - y = 3
x + 2y = 3
------
Modify for elimiination::
2x - y = 3
2x + 4y = 6
-----
Subtract and solve for "y":
5y = 3
y = 3/5
-=======
Solve for "x"::
x + 2y = 3
x + (6/5) = (15/5)
x = 9/5
------------
Cheers,
Stan H.